OpenStudy (anonymous):

Jess observed that 60% of the people at a mall on a particular day shopped for clothes. If 2500 people at the mall did not shop for clothes that day, the number of people who shopped for clothes that day was ______. (only put numeric values, no other symbols)

OpenStudy (anonymous):

hi @meanie2467 : its wrong to expect direct answer... every one here to help other to get to the answer.... if 60% of people at the mall shopped for clothes then how much didnot shopped for cloths

OpenStudy (anonymous):

40% didnt

jigglypuff314 (jigglypuff314):

so 0.4 (people) = 2500 therefore (people) = 2500/0.4

OpenStudy (anonymous):

Cool .... so let the total number of people who came for shopping Be X then 40 % of X = 2500 ie 40/100 * X = 2500

OpenStudy (anonymous):

ok soo can u just mulyiply 2500 by 60 and divide by 40

OpenStudy (anonymous):

x will be 6250 i think so .. bettr use calc.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

(2500*60)/40= 3750 .....is this right?